# Gradient-free stochastic optimization for additive models

### v2

@@ 4 Proofs
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable function such that $\left\lVert\nabla f(\boldsymbol{x})-\nabla f(\boldsymbol{x}^{\prime})\right\rVert\leq\bar{L}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$, where $\bar{L}>0$. Let Assumption 1 hold and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$. Then, for all $t\in[T]$ we have
@@
Note that $r_{t,i}$ and $-r_{t,i}$ have the same distribution. Therefore, $\mathbf{E}[G_{i}|\boldsymbol{x}_{t}]=0$ and we can write
@@
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $\bar{L},L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are updates of Algorithm 1 with $\Theta=\mathbb{R}^{d}$ and with the gradient estimators defined by (4). Then, for all $t\in[T]$ we have
@@
For $\alpha>0$ assume that $f$ is an $\alpha$-strongly convex function. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are the updates of Algorithm 1. Then, for all $t\in[T]$ and $\boldsymbol{x}\in\Theta$, we have
@@
Fix $\boldsymbol{x}\in\Theta$. Then, by the contraction property of the Euclidean projection we have $\left\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rVert^{2}\leq\left\lVert\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t}-\boldsymbol{x}\right\rVert^{2}$. Equivalently,
@@
$\eta_{t}=\begin{cases}\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}\quad\quad&\text{if $t<T_{0}$},\vspace{3mm}\\ \frac{4}{\alpha t}\quad\quad&\text{if $t\geq T_{0}$}.\end{cases}$
We consider separately the cases $T>T_{0}$ and $T\leq T_{0}$.
If $T>T_{0}$, then for $t\geq T_{0}$ we have $\eta_{t}=4/(\alpha t)$ and we can write
@@
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{t}\right)+\frac{d}{\alpha t}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha t}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{t}\left(\frac{\bar{L}}{\alpha tL^{2}}\right)^{\frac{1}{\beta}}\right).$
where $\texttt{A}_{3}=6\,\left((2\kappa_{\beta}^{2}/3)(\kappa\sigma^{2})^{\beta-1}\right)^{1/\beta}$ and $\texttt{A}_{4}=9\left(3\kappa\sigma^{2}/(2\kappa_{\beta}^{2})\right)^{\frac{1}{\beta}}/4$. Since $T\geq T_{0}$, applying [1, Lemma 32] leads to the bound
@@
$\delta_{T}$ $\leq\frac{2T_{0}}{T}\delta_{T_{0}+1}+\frac{d}{\alpha}\left(\frac{\beta\texttt{A}_{1}}{\beta+1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{\beta\texttt{A}_{2}}{2\beta+1}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
If $T_{0}=0$ then the proof is complete for the case $T>T_{0}$. Otherwise, for any $t\leq T_{0}$, we have $\eta_{t}=1/(18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa)$ and $h_{t}=(\frac{3\bar{L}}{\alpha T}\cdot\frac{\kappa\sigma^{2}}{2L^{2}\kappa_{\beta}^{2}})^{\frac{1}{2\beta}}$. Note that
$\frac{4}{(T_{0}+1)\alpha}\leq\eta_{t}\leq\frac{4}{T_{0}\alpha},$
and from (8) we deduce that, for all $t\leq T_{0}$,
$\delta_{t+1}\leq\delta_{t}\left(1-\frac{2}{T_{0}+1}\right)+\frac{2d}{\alpha T_{0}}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha}\right)^{\frac{\beta-1}{\beta}}\left(T^{-\frac{\beta-1}{\beta}}+\frac{T^{\frac{1}{\beta}}}{T_{0}}\right)+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T_{0}}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Note that $1-2/(T_{0}+1)\leq 1$. Thus, by taking into account the definition of $T_{0}$ and the fact that $T>T_{0}$ we get
$\frac{2T_{0}}{T}\delta_{t+1}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\frac{4d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Therefore,
$\delta_{T}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right),$
@@
where $\texttt{A}_{1}=\texttt{A}_{3}\left(\beta/(\beta+1)+8\right)$ and $\texttt{A}_{2}=\texttt{A}_{4}\left(\beta/(2\beta+1)+4\right)$.
Consider now the case $T\leq T_{0}$. Then from (8) we get that, for all $t\leq T$,
@@
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{T_{0}+1}\right)+\frac{2d}{\alpha T_{0}}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha}\right)^{\frac{\beta-1}{\beta}}\left(T^{-\frac{\beta-1}{\beta}}+\frac{T^{\frac{1}{\beta}}}{T_{0}}\right)+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T_{0}}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Thus,
$\delta_{T+1}$ $\leq\delta_{1}\left(1-\frac{2}{T_{0}+1}\right)^{T}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Since $(1-\lambda)^{T}\leq\exp(-\lambda T)\leq 1/(\lambda T)$ for all $T,\lambda>0$, we deduce from the previous display that
$\delta_{T+1}$ $\leq\frac{T_{0}+1}{2T}\delta_{1}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
By using the fact that $\lambda+1\leq 2\lambda$ for all $\lambda\geq 1$ we finally get
$\delta_{T+1}$ $\leq\frac{2T_{0}}{T+1}\delta_{1}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{2\bar{L}}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{2\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{2\bar{L}}{\alpha(T+1)L^{2}}\right)^{\frac{1}{\beta}}\right)$
@@
$\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha(T+1)}\delta_{1}+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{\bar{L}}{\alpha(T+1)L^{2}}\right)^{\frac{1}{\beta}}\right),$
where $\texttt{A}_{1}=2^{\frac{3\beta-1}{\beta}}\texttt{A}_{3}$ and $\texttt{A}_{2}=2^{\frac{2\beta+1}{\beta}}\texttt{A}_{4}$.
@@
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\leq\frac{a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right]}{2\eta_{t}}+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3\eta_{t}}{4}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)-\frac{\alpha}{4}a_{t}.$
Let $b_{t}=\mathbf{E}\left[\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}\right]$. Using the definition $\eta_{t}=4/(\alpha(t+1))$ and taking the expectation we obtain
$\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{\alpha}{4}\left(t\left(b_{t}-b_{t+1}\right)-b_{t}\right)+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3}{2\alpha(t+1)}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right).$
Summing up both sides of this inequality form $1$ to $T$ and using the fact that
$\sum_{t=1}^{T}t\left((t+1)\left(b_{t}-b_{t+1}\right)-b_{t}\right)\leq 0$
@@
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{d}{\alpha}\sum_{t=1}^{T}\left(t(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3t}{2(t+1)}\kappa\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)\right).$
Substituting here $h_{t}=(\frac{3}{2t}\frac{\kappa\sigma^{2}}{\kappa_{\beta}^{2}L^{2}})^{\frac{1}{2\beta}}$ yields
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{9G^{2}dT}{\alpha}+\frac{d}{\alpha}\sum_{t=1}^{T}\left(\texttt{A}_{3}(L^{2}t)^{\frac{1}{\beta}}+\texttt{A}_{4}d\bar{L}^{2}\left(L^{2}t\right)^{-\frac{1}{\beta}}\right),$
where $\texttt{A}_{3}=2(3\kappa\sigma^{2}/2)^{\frac{\beta-1}{\beta}}{\kappa}_{\beta}^{\frac{2}{\beta}}$ and $\texttt{A}_{4}=(1/2)\left(3/2\right)^{\frac{2\beta+1}{\beta}}\kappa^{\frac{\beta+1}{\beta}}\left(\sigma/\kappa_{\beta}\right)^{\frac{2}{\beta}}$. Using the inequality $\sum_{t=1}^{T}t^{-\frac{1}{\beta}}\leq(\beta/(\beta-1))T^{\frac{\beta-1}{\beta}}$ we find
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{9G^{2}dT}{\alpha}+\frac{d}{\alpha}\left(\texttt{A}_{3}L^{\frac{2}{\beta}}+\frac{\beta\texttt{A}_{4}}{\beta-1}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{\frac{\beta+1}{\beta}}.$
Dividing both sides of this inequality by $T(T+1)/2$ and applying Jensen’s inequality we finally get
$\mathbf{E}\left[f(\bar{\boldsymbol{x}}_{T})-f(\boldsymbol{x})\right]\leq\frac{18G^{2}d}{\alpha T}+\frac{d}{\alpha}\left(\texttt{A}_{1}L^{\frac{2}{\beta}}+\texttt{A}_{2}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}_{1}=2\texttt{A}_{3}$ and $\texttt{A}_{2}=2\beta\texttt{A}_{4}/(\beta-1)$.

### v3

@@ 4 Proofs
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable function such that $\left\lVert\nabla f(\boldsymbol{x})-\nabla f(\boldsymbol{x}^{\prime})\right\rVert\leq\bar{L}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$, where $\bar{L}>0$. Let Assumption 1 hold and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the realization of Algorithm 1 with $\Theta=\mathbb{R}^{d}$. Then, for all $t\in[T]$ we have
@@
Note that $r_{t,i}$ and $r_{t,i}$ have the same distribution. Therefore, $\mathbf{E}[G_{i}|\boldsymbol{x}_{t}]=0$ and we can write
@@
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $\bar{L},L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ is the realization of Algorithm 1 with $\Theta=\mathbb{R}^{d}$ and with the gradient estimators defined by (4). Then, for all $t\in[T]$ we have
@@
For $\alpha>0$ assume that $f$ is an $\alpha$-strongly convex function. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ is the realization of Algorithm 1. Then, for all $t\in[T]$ and $\boldsymbol{x}\in\Theta$, we have
@@
Fix $\boldsymbol{x}\in\Theta$. From the definition of the Euclidean projection we have $\left\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rVert^{2}\leq\left\lVert\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t}-\boldsymbol{x}\right\rVert^{2}$. Equivalently,
@@
$\eta_{t}=\begin{cases}\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}\quad\quad&\text{if $t{\leq}T_{0}$},\vskip 8.53581pt\\ \frac{4}{\alpha t}\quad\quad&\text{if $t\geq T_{0}+1$},\end{cases}$
and
$\frac{4}{(T_{0}+1)\alpha}\leq\eta_{t}\leq\frac{4}{T_{0}\alpha},\quad\text{for}\ t\leq T_{0}.$ (9)
We consider separately the cases $T\geq T_{0}+1$ and $T\leq T_{0}$.
If $T\geq T_{0}+1$ and $T\in\{1,2\}$ then, using (8), it is easy to check that the result of the theorem holds true.
Consider now the case $T\geq 3$ and $T\geq T_{0}+1$, that is, $T\geq T_{0}^{\prime}+1$, where $T_{0}^{\prime}=\max(T_{0},2)$. For any $t\geq T_{0}+1$ we have $\eta_{t}=4/(\alpha t)$ and we can write
@@
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{t}\right)+\texttt{A}\frac{d}{\alpha t}\left({\left(\frac{1}{\alpha t}\right)^{\frac{\beta-1}{\beta}}}+{\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{t}\left(\frac{1}{\alpha t}\right)^{\frac{1}{\beta}}}\right),$
where $\texttt{A}>0$ depends only on $L,\bar{L}$}, $\beta$, and $\sigma^{2}$, and its value may vary across different appearances throughout this proof. Since $T\geq{T_{0}^{\prime}+1\geq 3}$, applying [1, Lemma 32] leads to the bound
@@
$\delta_{T}$ $\leq\frac{2{T_{0}^{\prime}}}{T}\delta_{{T_{0}^{\prime}}+1}+\texttt{A}\frac{d}{\alpha}\left({\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}}+{\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}}\right).$ (10)
If $T_{0}\leq 2$ (that is, $T_{0}^{\prime}=2$) we use (8) with $t=1,t=2,$ the definition of $T_{0}$, (9), and the definition of $h_{t}$ to obtain
@@
$\delta_{{T_{0}^{\prime}}+1}=\delta_{3}\leq\delta_{1}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha}\right)^{\frac{\beta-1}{\beta}}+{d}\left(\frac{1}{\alpha}\right)^{\frac{1}{\beta}}\right).$
Combining this bound with (10) completes the proof for the case $_{0}^{\prime}=2, TT$$\geq T_{0}^{\prime}+1$.
Let now $T_{0}^{\prime}\geq 3$ (which implies $T_{0}=T_{0}^{\prime}$), and $T\geq T_{0}^{\prime}+1$. From (8), (9) and the definition of $h_{t}$ we obtain that, for any $t\leq T_{0}$,
@@
$\delta_{t+1}\leq\delta_{t}\left(1-\frac{2}{T_{0}+1}\right)+\texttt{A}\frac{d}{\alpha T_{0}}\left(\left(\frac{1}{\alpha}\right)^{\frac{\beta-1}{\beta}}\left(T^{-\frac{\beta-1}{\beta}}+\frac{T^{\frac{1}{\beta}}}{T_{0}}\right)+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T_{0}}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right).$ (11)
By iterating (11) with the simplification $1-2/(T_{0}+1)\leq 1$, and taking into account the definition of $T_{0}$ and the inequality $T\geq T_{0}+1$ we find that
@@
$\frac{2T_{0}}{T}\delta_{{T_{0}}+1}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\texttt{A}\frac{{}d}{\alpha}\left(\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right).$
This bound and (10) imply that
$\delta_{T}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right).$
The proof for the case $T\geq T_{0}+1$ is complete.
Consider now the case $T\leq T_{0}$. In this case (11) holds for all $t\leq T$ and we have the bound
$\delta_{T+1}$ $\leq\delta_{1}\left(1-\frac{2}{T_{0}+1}\right)^{T}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right).$
Since $(1-\lambda)^{T}\leq\exp(-\lambda T)\leq 1/(\lambda T)$ for all $T>0,\lambda\in(0,1)$ we obtain
$\delta_{T+1}$ $\leq\frac{T_{0}+1}{2T}\delta_{1}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{1}{\alpha T}\right)^{\frac{1}{\beta}}\right).$
By using the fact that $4TT_{0}\geq(T+1)(T_{0}+1)$ we finally get
$\delta_{T+1}$ $\leq\frac{2T_{0}}{T+1}\delta_{1}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{1}{\alpha(T+1)}\right)^{\frac{1}{\beta}}\right)$
$\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha(T+1)}\delta_{1}+\texttt{A}\frac{d}{\alpha}\left(\left(\frac{1}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{1}{\alpha(T+1)}\right)^{\frac{1}{\beta}}\right).$
@@
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})$ $\leq\frac{a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right]}{2\eta_{t}}+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+$
$\quad+\frac{3\eta_{t}}{4}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)-\frac{\alpha a_{t}}{4}.$
Let $b_{t}=\mathbf{E}\left[\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}\right]$. Substituting here $\eta_{t}=4/(\alpha(t+1))$ and taking the expectation we obtain
$\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]$ $\leq\frac{\alpha}{8}\left((t+1)\left(b_{t}-b_{t+1}\right)-2b_{t}\right)+$
$\quad+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3}{\alpha(t+1)}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right).$
Summing up both sides of this inequality from $1$ to $T$ and using the fact that
$\sum_{t=1}^{T}t\left((t+1)\left(b_{t}-b_{t+1}\right)-2b_{t}\right)\leq 0,$
@@
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq$
$\qquad\leq\frac{d}{\alpha}\sum_{t=1}^{T}\left(t(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3t}{t+1}\kappa\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)\right).$
Substituting here $h_{t}=(\frac{3}{2t}\frac{\kappa\sigma^{2}}{\kappa_{\beta}^{2}L^{2}})^{\frac{1}{2\beta}}$ we get the inequality
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]$ $\leq\frac{18G^{2}d\kappa T}{\alpha}+\texttt{A}\frac{d}{\alpha}\sum_{t=1}^{T}\left(t^{\frac{1}{\beta}}+dt^{-\frac{1}{\beta}}\right)\leq$
$\quad\leq\frac{18G^{2}d\kappa T}{\alpha}+\texttt{A}^{\prime}\frac{d}{\alpha}\left(1+dT^{-\frac{2}{\beta}}\right)T^{\frac{\beta+1}{\beta}}.$
Here $\texttt{A},\texttt{A}^{\prime}$ are positive constants depending only on $L$, $\bar{L}$, $\beta$, and $\sigma^{2}$. Dividing both sides of this inequality by $T(T+1)/2$ and applying Jensen’s inequality we get the result of the theorem.